A committee of 7 is chosen from 9 men and 8 women. The number of committees that contain at least 6 men is: A 708 B 602 C 954 D 552
Answer #1

Committee should contain 7 persons.
To find: Number of committees with atleast 6 men.

Possibilities:

1) 6 men + 1 woman
2) 7 men + 0 woman

case 1: 6 men + 1 woman:

Number of commities= (9C6)*(8C1)=84*8=672

case 2: 7 men + 0 woman:

Number of commities=(9C7)*(8C0)=36*1=36

Total number of committees with atleast 6 men =672+36=708.

Therefore Answer is A) 708